# MMDU: A Multi-Turn Multi-Image Dialog Understanding Benchmark and Instruction-Tuning Dataset for LVLMs

### v1

@@ Abstract
Generating natural and meaningful responses to communicate with multi-modal human inputs is a fundamental capability of Large Vision-Language Models (LVLMs). While current open-source LVLMs demonstrate promising performance in simplified scenarios such as single-turn single-image input, they fall short in real-world conversation scenarios such as following instructions in a long context history with multi-turn and multi-images. Existing LVLM benchmarks primarily focus on single-choice questions or short-form responses, which do not adequately assess the capabilities of LVLMs in real-world human-AI interaction applications. Therefore, we introduce MMDU, a comprehensive benchmark, and MMDU-45k, a large-scale instruction tuning dataset, designed to evaluate and improve LVLMs’ abilities in multi-turn and multi-image conversations. We employ the clustering algorithm to find the relevant images and textual descriptions from the open-source Wikipedia and construct the question-answer pairs by human annotators with the assistance of the GPT-4o model. MMDU has a maximum of 18k image+text tokens, 20 images, and 27 turns, which is at least 5$\times$ longer than previous benchmarks and poses challenges to current LVLMs. Our in-depth analysis of 15 representative LVLMs using MMDU reveals that open-source LVLMs lag behind closed-source counterparts due to limited conversational instruction tuning data. We demonstrate that fine-tuning open-source LVLMs on MMDU-45k significantly address this gap, generating longer and more accurate conversations, and improving scores on MMDU and existing benchmarks (MMStar: +1.1%, MathVista: +1.5%, ChartQA: +1.2%). Our contributions pave the way for bridging the gap between current LVLM models and real-world application demands. This project is available at https://github.com/Liuziyu77/MMDU.
@@ 2.2 Benchmark Construction
In the process of constructing the dataset, we implemented two stringent measures to ensure its quality: (1) We combined automated and manual screening methods to select images and texts that meet our standards. Specifically, we performed an initial screening using clustering techniques on a large-scale image and text database, automatically removing low-quality, blurry, or irrelevant images and texts. This ensured that the image combinations and their corresponding texts were of high quality and relevance. (2) To avoid hallucinations and errors in the model-generated dialogues, we enforced strict quality control on the texts generated by GPT-4o. We introduced a multi-round manual review mechanism. Each set of Q&A underwent at least two rounds of manual review: the first round involved preliminary checks by regular reviewers, and the second round involved in-depth examination and modification by experts. This ensured that the final dataset was not only accurate but also of high academic and practical value.
@@ 2.3 Evaluation
Inspired by NLP research that leverages strong LLMs as judges [6], we have developed an evaluation pipeline using GPT-4o to evaluate model performance. Specifically, following the generation of model predictions on our benchmark dataset, GPT-4o evaluates these predictions across various dimensions for each turn and sample, comparing them against reference answers. The aggregated results across multiple turns are averaged to derive sample scores, and the average scores across all samples constitute our benchmark scores.
To ensure a comprehensive and nuanced evaluation, we have identified six dimensions: Creativity, Richness, Visual Perception, Logical Coherence, Answer Accuracy, and Image Relationship Understanding. To guide GPT-4o in providing balanced and equitable assessments, we have meticulously crafted evaluation prompts for each dimension. Each dimension’s score range of 10 is divided into five intervals (0-2, 2-4…8-10), with corresponding judgment criteria established for each interval. GPT-4o follows these criteria to conduct judgment processes and deliver final scores for each dimension. As illustrated in Fig 3, guided by our prompts, GPT-4o assesses the assistant’s responses against reference answers, offering both a reasonable score and a transparent judgment process. Please refer to the supplementary material B for our judgment prompts.
@@ 3.1 Dataset Construction
3.2 Dataset Statistics
@@ 4 Experiments
We report the performance of four closed-source API models: QWen-VL-Max [13], Claude3 [14], GPT-4-turbo [1] and GPT-4o [15]. We also present the performance of 11 LVLMs including Monkey [16], Idefics2 [17], LLaVa1.5 7B/13B [18], Deepseek-VL [19], MiniCPM-v-2.5 [20, 21], Yi-VL [22], InternVL-Chat-V1.5 [23], InternLM-XC2 [7], Qwen-VL-7B [13] and LLaVa1.6 [24]. Please refer to the supplementary material D for the details of our baselines.
@@ 4.1 Main Results on MMDU
Table 2 presents the benchmarking results on our MMDU benchmark. Our key findings are summarized as follows. (1) Our benchmark poses significant challenges to current LVLMs. Notably, even the advanced GPT-4o model achieves an average accuracy of only 70.2%, while open-source LVLMs achieve merely 42.8% or lower, indicating substantial room for improvement. (2) We observe a significant performance gap between closed-source LVLMs and open-source LVLMs. We speculate that this disparity arises from the scarcity of open-source instruction tuning data with multi-turn and multi-image capabilities, leading to limited improvement in open-source LVLMs. This inspired us to collect and release MMDU-45k, a valuable resource for the open-source community, to bridge this gap.
@@ 5 Related Work
The rapid advancements in Large Vision-Language Models (LVLMs)[13, 14, 1, 15, 16, 17, 18, 19, 20, 21, 22, 23, 7, 13, 24, 30, 31] have spurred the development of comprehensive evaluation benchmarks to assess their capabilities across various tasks and domains. Numerous benchmarks [32, 26, 33, 34, 35, 36, 37, 4, 8] aim to provide a standardized and objective way to measure the performance of LVLMs and track their progress toward achieving general multi-modal understanding and reasoning.
@@ 6 Conclusion
Supplementary Material
In this supplementary material, we offer further details regarding the proposed MMDU and MMDU-45k, along with additional experimental discussions aimed at comprehensive benchmarking. Specifically, Section A includes our project URL and benchmark download URL. Sections B and C delve into the specifics of MMDU and MMDU-45k, respectively. Our evaluation section D provides in-depth analysis and discussion. Section E provides the datasheets for MMDU and MMDU-45k.
@@ C.2 The Richness and Diversity of MMDU-45k
Appendix D More Evaluation Details
@@ D.2 More cases
To clarify the testing and evaluation process of MMDU, we display three question-answer pairs from MMDU in Fig. 14, 15 and 16. Due to space limitations, we cannot show a complete multi-turn, multi-image conversation, so we selected one question-answer pair for demonstration. Each case includes the relevant images and the ground truth. We also list the results of InternLM-Xcomposer2 and InternLM-Xcomposer2+MMDU-45k, showcasing the effectiveness of MMDU-45k in improving the model’s ability to handle multi-image, multi-turn conversations. Additionally, we provide the scoring results using GPT-4o, including the reasons and given scores.

### v2

@@ Abstract
Generating natural and meaningful responses to communicate with multi-modal human inputs is a fundamental capability of Large Vision-Language Models (LVLMs). While current open-source LVLMs demonstrate promising performance in simplified scenarios such as single-turn single-image input, they fall short in real-world conversation scenarios such as following instructions in a long context history with multi-turn and multi-images. Existing LVLM benchmarks primarily focus on single-choice questions or short-form responses, which do not adequately assess the capabilities of LVLMs in real-world human-AI interaction applications. Therefore, we introduce MMDU, a comprehensive benchmark, and MMDU-45k, a large-scale instruction tuning dataset, designed to evaluate and improve LVLMs’ abilities in multi-turn and multi-image conversations. We employ the clustering algorithm to find the relevant images and textual descriptions from the open-source Wikipedia and construct the question-answer pairs by human annotators with the assistance of the GPT-4o model. MMDU has a maximum of 18k image+text tokens, 20 images, and 27 turns, which is at least 5$\times$ longer than previous benchmarks and poses challenges to current LVLMs. Our in-depth analysis of 15 representative LVLMs using MMDU reveals that open-source LVLMs lag behind closed-source counterparts due to limited conversational instruction tuning data. We demonstrate that fine-tuning open-source LVLMs on MMDU-45k significantly addresses this gap, generating longer and more accurate conversations, and improving scores on MMDU and existing benchmarks (MMStar: +1.1%, MathVista: +1.5%, ChartQA: +1.2%). Our contributions pave the way for bridging the gap between current LVLM models and real-world application demands. This project is available at https://github.com/Liuziyu77/MMDU.
@@ 2.2 Benchmark Construction
In the process of constructing the dataset, we implemented two stringent measures to ensure its quality: (1) We combined automated and manual screening methods to select images and texts that meet our standards. Specifically, we performed an initial screening using clustering techniques on a large-scale image and text database, automatically removing low-quality, blurry, or irrelevant images and texts. This ensured that the image combinations and their corresponding texts were of high quality and relevance. (2) To avoid hallucinations and errors in the model-generated dialogues, we enforced strict quality control on the texts generated by GPT-4o. We introduced a multi-round manual review mechanism. Each set of Q&A underwent at least two rounds of manual review: the first round involved preliminary checks by regular reviewers, and the second round involved in-depth examination and modification by experts.
For the two rounds of manual review, our experts reviewed and corrected (by removing or rewriting) any hallucinations and errors to ensure that all dialogues are accurate. To facilitate verification, we designed a specialized web UI that allows for quick browsing and modification of data content. Please refer to the Appendix  B.5 to check the interface of our web UI used for the human check process.
During the data annotation process, all our annotators were either junior PhD-level students or senior researchers, with a total of 20 participants. Senior researchers or PhD students with relevant professional backgrounds were selected as experts. Before data annotation, all annotators underwent training, and a small subset of data was pre-annotated. Once the pre-annotation results met the required standards, the subsequent annotation process was carried out.
Since our images are sourced from Wikipedia entries, each image is accompanied by a corresponding caption and all related information from the Wiki entry where the image is found. During the manual annotation process, annotators can easily understand the specific content and background information of each image by reading the Wiki entry. Therefore, there is no risk of introducing extra annotation errors due to misunderstanding of the image content. The various strategies mentioned above ensured that the final dataset was not only accurate but also of high academic and practical value.
@@ 2.3 Evaluation
Evaluating the subjective, open-ended, free-form, and long-context visual question-answering pairs is indeed challenging. Traditional metrics (e.g., BLEU-4, CIDEr) often suffer from shortcomings like neglecting semantic understanding and failing to capture long-distance dependencies, they are not popular choices in recent days.
Inspired by NLP research that leverages strong LLMs as judges [6], we have developed an evaluation pipeline using GPT-4o to evaluate model performance. Specifically, following the generation of model predictions on our benchmark dataset, GPT-4o evaluates these predictions across various dimensions for each turn and sample, comparing them against reference answers. The aggregated results across multiple turns are averaged to derive sample scores, and the average scores across all samples constitute our benchmark scores. This method excels at understanding context and semantics, providing more accurate evaluations of visual content, and capturing long-distance dependencies, which traditional metrics often miss.
To ensure a comprehensive and nuanced evaluation, we have identified six dimensions: Creativity, Richness, Visual Perception, Logical Coherence, Answer Accuracy, and Image Relationship Understanding. To guide GPT-4o in providing balanced and equitable assessments, we have meticulously crafted evaluation prompts for each dimension. Each dimension’s score range of 10 is divided into five intervals (0-2, 2-4…8-10), with corresponding judgment criteria established for each interval. GPT-4o follows these criteria to conduct judgment processes and deliver final scores for each dimension. As illustrated in Fig 3, guided by our prompts, GPT-4o assesses the assistant’s responses against reference answers, offering both a reasonable score and a transparent judgment process. Please refer to the Appendix B for our judgment prompts.
@@ 3.1 Dataset Construction
In the manual data inspection phase, due to the large volume of data in the MMDU-45k, it was not feasible to review all of the data, so we sampled $5\%$ of the dataset for inspection. Statistical analysis showed that the probability of hallucinations and errors in this subset was less than $5\%$, indicating a high level of reliability.
@@ 4 Experiments
We report the performance of four closed-source API models: QWen-VL-Max [13], Claude3 [14], GPT-4-turbo [1] and GPT-4o [15]. We also present the performance of 11 LVLMs including Monkey [16], Idefics2 [17], LLaVa1.5 7B/13B [18], Deepseek-VL [19], MiniCPM-v-2.5 [20, 21], Yi-VL [22], InternVL-Chat-V1.5 [23], InternLM-XC2 [7], Qwen-VL-7B [13] and LLaVa1.6 [24]. Please refer to the Appendix D for the details of our baselines.
@@ 4.1 Main Results on MMDU
Tab. 2 presents the benchmarking results on our MMDU benchmark. Our key findings are summarized as follows. (1) Our benchmark poses significant challenges to current LVLMs. Notably, even the advanced GPT-4o model achieves an average accuracy of only 70.2%, while open-source LVLMs achieve merely 42.8% or lower, indicating substantial room for improvement. (2) We observe a significant performance gap between closed-source LVLMs and open-source LVLMs. We speculate that this disparity arises from the scarcity of open-source instruction tuning data with multi-turn and multi-image capabilities, leading to limited improvement in open-source LVLMs. This inspired us to collect and release MMDU-45k, a valuable resource for the open-source community, to bridge this gap.
For Tab. 2, we found that InternLM-Xcomposer2 benefits more from MMDU than LLaVA1.5. This is because, as a more recent model, InternLM-Xcomposer2 uses a different LLM backbone, a more powerful architecture, larger pre-training and SFT data, and more advanced training strategies compared to LLaVa1.5. These advantages may provide InternLM-Xcomposer2 with stronger generalization capabilities.
@@ 4.2 Fine-tuning Results using MMDU-45k
4.3 Multi-Image Benchmark Results
Additionally, we test the model finetuned with MMDU-45k on several multi-image benchmarks. We evaluate five benchmarks: MMMU [27], BLINK [30], Qbench2 [31], Mantis [32], and MMDU. For MMMU, only the results of multi-image questions are considered, and for Mantis, tests are conducted using both the "merge" and "sequence" methods. As shown in Tab. 4, LLaVa1.5+MMDU-45k achieved significant improvements across all multi-image benchmarks, with the most notable improvement observed in Mantis (sequence), reaching a $6.9\%$ increase. This indicates that MMDU-45k greatly aids in enhancing the model’s multi-image understanding capabilities, significantly addressing the model’s shortcomings in reasoning within multi-image scenarios due to a lack of multi-image pre-train data.
Models MMMU BLINK Qbench2 Mantis Mantis MMDU
(multi-pics) (sequence) (merge)
LLaVa-1.5 27.7 37.1 46.0 37.8 41.9 32.2
+MMDU-45k 29.8 40.1 48.5 44.7 44.7 37.2
$\Delta$ +2.1 +3.0 +2.5 +6.9 +2.8 +5.0
Table 4: Multi-image benchmark test results. We evaluated five multi-image benchmarks: MMMU [27], BLINK [30], Qbench2 [31], Mantis [32], and MMDU.
4.4 Ablation Study on Token Length and SFT Strategies
We test the LLaVa1.5 baseline (context-window length is 2-4k) and LLaVa1.5 (SFT on MMDU-45k, we extend the context-window length to 8k with the RoPE interpolation) model with different context lengths. From Tab. 5, we observe that: (1) As the context length of the model increases, the performance also improves. (2) Finetuning on MMDU can increase the context window size of the LLaVA model.
Additionally, we compare different SFT strategies for training the model, including "Continue training" and "Add to the existing pool." The results in Tab. 5 show that the final outcomes achieved by both methods are essentially the same.
Table 5: Ablation Study on Token Length and SFT Strategies
Models Max C R VP LC AA IRU Overall
tokens Score
LLaVa-1.5 2k 19.0 19.0 21.8 29.3 22.5 19.6 20.9
4k 25.4 25.6 31.1 40.8 32.9 29.5 30.0
LLaVa-1.5+MMDU-45k 2k 20.0 20.1 22.1 29.4 23.5 21.6 22.3
4k 31.5 32.3 34.9 45.0 36.3 33.8 34.9
8k 34.2 34.3 36.1 48.2 39.6 34.7 37.1
Continue training - 34.3 34.5 36.7 47.2 38.5 35.5 37.2
Add to the existing pool - 34.3 36.3 37.1 47.3 38.9 35.7 37.3
@@ 5 Related Work
The rapid advancements in Large Vision-Language Models (LVLMs)[13, 14, 1, 15, 16, 17, 18, 19, 20, 21, 22, 23, 7, 13, 24, 33, 34, 35, 36] have spurred the development of comprehensive evaluation benchmarks to assess their capabilities across various tasks and domains. Numerous benchmarks [37, 26, 38, 39, 40, 41, 42, 4, 8] aim to provide a standardized and objective way to measure the performance of LVLMs and track their progress toward achieving general multi-modal understanding and reasoning.
@@ 6 Conclusion
Acknowledgments
This project is funded in part by Shanghai Artificial lntelligence Laboratory, the National Key R&\&&D Program of China (2022ZD0160201), the Centre for Perceptual and Interactive Intelligence (CPII) Ltd under the Innovation and Technology Commission (ITC)’s InnoHK. Dahua Lin is a PI of CPII under the InnoHK.
Appendices
In this appendix, we offer further details regarding the proposed MMDU and MMDU-45k, along with additional experimental discussions aimed at comprehensive benchmarking. Specifically, Appendix A includes our project URL and benchmark download URL. Appendix B and C delve into the specifics of MMDU and MMDU-45k, respectively. Our evaluation Appendix D provides in-depth analysis and discussion. Appendix E provides the datasheets for MMDU and MMDU-45k.
@@ B.4 Related Work with Existing Benchmarks
B.5 More Details about Human Annotators
We designed a dedicated data manipulation Web UI for manual data inspection, and its interface is shown in Fig. 11.
@@ C.2 The Richness and Diversity of MMDU-45k
Appendix D More Details
@@ D.2 More cases
To clarify the testing and evaluation process of MMDU, we display three question-answer pairs from MMDU in Fig. 15, 16 and 17. Due to space limitations, we cannot show a complete multi-turn, multi-image conversation, so we selected one question-answer pair for demonstration. Each case includes the relevant images and the ground truth. We also list the results of InternLM-Xcomposer2 and InternLM-Xcomposer2+MMDU-45k, showcasing the effectiveness of MMDU-45k in improving the model’s ability to handle multi-image, multi-turn conversations. Additionally, we provide the scoring results using GPT-4o, including the reasons and given scores.
@@ D.3 Finetune details
D.4 Cluster Accuracy
In the process of constructing multi-item clusters, we ensure clustering accuracy through two key steps: We first utilize the inherent tags or labels associated with each wiki item for clustering. These tags and labels are manually annotated in wiki items, providing a high level of accuracy. Additionally, we further employ image captions for clustering, ensuring that the resulting clusters are largely free of noise and errors.
To verify the accuracy of the clustering, we conducted the following experiment: We randomly sampled five sets of data from the MMDU-45k dataset, each containing 100 entries. The images from each entry were input into the GPT-4o model, which was tasked with evaluating whether the multiple images were related and could be grouped into the same cluster. The resulting accuracy rates were recorded as follows: $94\%$, $90\%$, $92\%$, $89\%$, $91\%$. This indicates that the clustering accuracy of MMDU is very high. The prompt is shown in Fig. 18.
D.5 Reasoning in MMDU
When answering questions from MMDU, the model must process interleaved image and text inputs while simultaneously understanding the relationships between multiple images which requires various visual reasoning capabilities. We believe that MMDU’s visual reasoning capabilities can be reflected in the following aspects:
Visual Perception Reasoning: to interpret and infer information from single or multiple images, recognizing relevant details, patterns, and relationships between the images.
Visual Relationship Reasoning: to understand and articulate the connections and relationships between different images, determining how they interact or contribute to the overall dialogue.
High-Level Semantic Reasoning: to grasp the deeper meanings, symbols, or abstract concepts represented across multiple images.
Cross-modal reasoning: to extract visual information from an image and combine it with textual or conversational cues to infer knowledge or information that goes beyond the content of the image.
These reasoning abilities are also considered when constructing the QA pairs for MMDU. In Fig. 19 and Fig. 20, we’ve included cases that illustrate the dataset’s visual reasoning capabilities.
@@ D.5 Reasoning in MMDU
Figure 18: Prompt for Clustering Accuracy Verification
Figure 19: Left: Visual Perception Reasoning. The most basic visual reasoning. Right: Visual Relationship Reasoning. Understand and articulate the connections and relationships between different images.
Figure 20: Left: High-Level Semantic Reasoning. Grasping the deeper meanings, symbols, or abstract concepts across multiple images. Right: Cross-modal reasoning. Extracting visual information from an image and combining it with textual cues to infer knowledge or information that goes beyond the content of the image.
@@ E.7 Maintenance
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state our contributions, and the claims match the experimental results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations are discussed in Section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Detailed instructions for replicating the results are provided in Appendix A. Additionally, the code, model checkpoint, and data will be publicly released.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The data source link is presented in Appendix A. Detailed information to reproduce all experimental results is provided in Appendix D.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The experimental setting is presented in Section 4 and Appendix D.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Error bars are not reported because it would be too computationally expensive.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [No]
  39. Justification: The experiment did not require too much time.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research adheres to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The potential impacts are discussed in Section 6.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: We will require users to adhere to specific usage guidelines to access the model and datasets, ensuring that they are used responsibly and to mitigate the risk of misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have properly credited the creators or original owners of assets used in the paper and we use the license CC-BY 4.0.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: All new assets introduced in this paper will be well documented upon their release.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.